# Towards Comprehensive and Efficient Post Safety Alignment of Large Language Models via Safety Patching

### v1

@@ 4 Experiments
4.1 Dataset and Evaluation Metrics
Safety Benchmark We evaluate the safety enhancement of the target backbone with PSA under two settings: the vanilla harmful prompt and the jailbreak attack, respectively. For the former, our evaluation is performed on the AdvBench (Zou et al., 2023) and Beavertails (Ji et al., 2023) datasets. The harmful data $D_{h}$ used for training comes from a subset of the Beavertails training set. We sample up to 100 sample from each of the 14 safety categories it contains. We adopt the harmful rate as the evaluation metric and it is automatically calculated by a Longformer-based (Beltagy et al., 2020) classifier provided by Wang et al. (2023). It exhibits identical performance as GPT-4 and human annotators to judge whether a response is harmful or not. Please refer to Appendix C for more details of the classifier. For the jailbreak attack, it is evaluated on AdvBench only. The harmful data now comes from the portion induced by the jailbreak attack method ReNeLLM (Ding et al., 2023). After applying SafePatching, we re-test the resulting PSA model’s performance under the ReNeLLM attack. Following (Zou et al., 2023; Qi et al., 2023), we adopt ASR (Attack Success Rate) as the evaluation metric. And ASR is also automatically derived from the above classifier.
Over-Safety Benchmark The evaluation for over-safety mitigation is performed on XSTest (Röttger et al., 2023) and OKTest (Shi et al., 2024). XSTest comprises 250 safe prompts across ten prompt types that well-calibrated models should not refuse to comply with. And in OKTest, there are 300 test samples characterized by safe questions containing harmful words. As suggested by Röttger et al. (2023), we hire three human annotators for manual evaluation of refusal rate.
Utility Benchmark To evaluate the utility of the backbone model, following Touvron et al. (2023b), we conduct evaluations across four crucial dimensions: (1) World Knowledge: MMLU (5-shot) (Hendrycks et al., 2020); (2) Reasoning: 0-shot for HellaSwag (Zellers et al., 2019), ARC-challenge (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021) and BBH (3-shot) Suzgun et al. (2023); (3) MATH: GSM8K (8-shot) (Cobbe et al., 2021) and MATH (4-shot) (Hendrycks et al., 2021) and (4) Multi-Turn Instruction-Following: MT-bench (Zheng et al., 2024).
Figure 3: Results of SafePatching and baselines under the jailbreak attack of ReNeLLM (Ding et al., 2023). ASR is reported on the AdvBench.
@@ 4.2 Baselines and Comparison Methods
We evaluate SafePatching against the following PSA baselines, which could be divided into two categories: (1) Unlearning-based: GA (Yao et al., 2023), GA + Mismatch (Yao et al., 2023), Task Vector (Zhang et al., 2023b) and NPO (Zhang et al., 2024a); (2) Decoding-based: SafeDecoding (Xu et al., 2024), ROSE (Zhong et al., 2024) and Self-CD (Shi et al., 2024).
To verify the advantages of the controllable patching in conflict mitigation, we also compare SafePatching with the state-of-the-art model merging methods: (1) Average Merging (Wortsman et al., 2022), (2) Task Arithmetic (Ilharco et al., 2022a), (3) Fisher Merging (Matena and Raffel, 2022) and (4) TIES-Merging (Yadav et al., 2023).
Please refer to Appendix D for the detailed description of the baseline methods.
@@ 4.3 Implementation Details
SafePatching is a model-agnostic PSA framework that is compatible with any transformer-based LLM. In our experiments, all methods are implemented on LLaMA-2-Chat-7B and LLaMA-2-Chat-13B backbones with PyTorch (Paszke et al., 2019) and Transformer library (Wolf et al., 2020). For more detailed settings, please refer to the Appendix E.
Table 1: The overall results on the safety, over-safety and utility benchmarks with LLaMA-2-Chat-7B backbone. The evaluation metrics for safety and over-safety is harmful rate and refusal rate, respectively. We report the average results on all the utility dataset except for MT-bench (1 - 10).
@@
AdvBench Beavertails XSTest OKTest AVG. MT-Bench
Original 31.73 19.64 8 4.67 40.35 6.01
GA (Yao et al., 2023) 0.58 3.65 44.40 34.67 39.32 3.11
GA + Mismatch (Yao et al., 2023) 2.31 3.98 23.59 13.67 39.05 4.19
Task Vector (Zhang et al., 2023b) 1.54 5.39 68 53.67 36.57 4.73
NPO (Zhang et al., 2024a) 26.73 17.07 8.79 8.67 39.06 5.79
SafeDecoding (Xu et al., 2024) 0 0.01 80.80 59.67 - 5.72
Self-CD (Shi et al., 2024) 0 0.03 22.80 41.67 - 3.98
ROSE (Zhong et al., 2024) 0 0.02 43.20 40.33 - 4.14
SafePatching (Ours) 3.27 7.87 5.60 2.67 39.99 6.16
5 Results and Analysis
Table 1 and Table 7 (in Appendix F.2) demonstrates the performance comparison of SafePatching and recent unlearning-based and decoding-based baselines on different benchmarks.
@@ 5.1 Overall Results
SafePatching could effectively achieve safety enhancement, over-safety mitigation and utility preservation simultaneously. As shown in Table 1, while both unlearning-based and decoding-based methods can greatly improve the safety, they also lead to a notable increase in over-safety issues. When it comes to maintaining utility, baseline methods, especially gradient ascent-based unlearning methods (GA, GA + Mismatch and NPO), exhibit a significant decline in performance on the MT-bench, which assesses the model’s multi-turn conversational and instruction-following abilities. We do not report
Figure 4: ATGR of different methods to manifest the inference cost.
ATGR
SafeDecoding 1.03 $\times$
Self-CD 2.48 $\times$
ROSE 2.08 $\times$
SafePatching 1 $\times$
the AVG. results of these decoding-based methods due to their enormous inference costs, which make the entire testing process highly inefficient. Their loss of the utility can also be reflected in the significant performance degradation on MT-bench. The results shown in Figure 3 under the jailbreak attack setting exhibit a similar trend, indicating that SafePatching could also serve as an effective post-hoc remedy against jailbreak attacks.
Moreover, SafePatching is more efficient than both unlearning-based and decoding-based methods, as it does not require extra training data and inference cost. In Table 4, we showcase the inference speeds of various methods, measured using the average token generation time ratio (ATGR) metric (Xu et al., 2024). These decoding-based methods clearly slow down the model’s inference speed because of the extra operations on token probabilities during decoding. In contrast, our SafePatching approach maintains the original architecture and inference process, ensuring no additional overhead and eliminating the need for further adjustments during model deployment. Please refer to Appendix F.3 for detailed analysis on the training and inference costs for each method.
@@
SafePatching could even enhance utility in the PSA process. Surprisingly, SafePatching outperforms the original model on the MT-bench. We explore the reasons behind this improvement by analyzing GPT-4’s evaluation of the generated responses. We discover that responses from SafePatching more frequently address the user’s input directly, unlike the original model, which may criticize the phrasing of user queries due to over-safety issues (all input prompts in MT-bench are safe). In other words, SafePatching further optimizes the balance between being helpful and being harmless. Thus, responses from SafePatching are preferred by the GPT-4 evaluator, which is also aligned with human annotators’ preference for models to provide direct responses to safe and harmless questions. Detailed analysis can be found in Appendix G.3.
Table 2: The overall ablation results to verify the effectiveness of different components in SafePatching. We compare it with the SOTA model merging methods.
@@
AdvBench Beavertails XSTest OKTest AVG. MT-Bench
Original 31.73 19.64 8 4.67 40.35 6.01
Average Merging (Wortsman et al., 2022) 48.27 26.01 0.40 0 40.17 6.19
Task Arithmetic (Ilharco et al., 2022b) 44.42 21.21 2 3 40.46 6.19
Fisher Merging (Matena and Raffel, 2022) 88.65 62.72 0 0 38.49 6.05
TIES-Merging (Yadav et al., 2023) 18.27 16.98 2.40 3.50 40.23 6.04
w/ Intersect. Patch 32.50 19.88 6 5.33 40.42 5.98
w/o Rand. Retention 3.50 5.0 11.60 7.0 39.06 5.64
w/ Safety Patch 0 0.08 98.40 97.50 39.27 5.31
w/ Over-Safety Patch 87.69 65.45 0 0 37.21 5.97
SafePatching 3.27 7.87 5.60 2.67 39.99 6.16
5.2 Deeper Analysis on SafePatching
We further conduct extensive ablation studies to demonstrate the effectiveness of different components of SafePatching. The overall results are shown in Table 2.
Safety and over-safety patches are highly effective in achieving their specific goals. Whether incorporating safety patch or over-safety patch alone (denoted as “w/ Safety Patch” and “w/ Over-Safety Patch” in Table 2), the resulting model consistently excels in meeting the respective PSA objectives. This highlights the logic and effectiveness of using harmful data for gradient optimization to enhance safety and reduce over-safety independently.
Random retention of parameters in patches is essential for maintaining the utility of the backbone model. When we remove the operation of random retention (denoted as “w/o Rand. Retention” in Table 2), the utility of the resulting model decreases, revealing the conflict between the parameters of these two safety patches and the backbone model. Additionally, in the Appendix F.4, we examine how varying overall retention rates affect both the utility of the base model and the preservation of the patch’s unique attributes.
Our proposed Controllable Patching could effectively alleviate the conflict between the safety and over-safety patches. To demonstrate advantages of our SafePatching in mitigating conflicts between safety and over-safety patches, we compare it with the SOTA model merging methods. Note that these baselines are also integrated with the same random retention techniques (Yu et al., 2023). Although TIES-Merging is the most effective method for mitigating conflicts in model merging, it struggles to find a balance between safety and over-safety. This issue arises from its lack of controlled merging in areas specifically related to safety and over-safety, as demonstrated by SafePatching. More importantly, while these model merging methods succeed to increase the backbones’ utility, this enhancement has significantly compromised their safety. This trade-off is clearly unacceptable for delivering safe and reliable LLM services. We hope this trade-off serves as a reminder that future work on model merging should prioritize both improving utility and maintaining safety.
Additionally, recall that our merging of safety and over-safety patches occurs in the difference set of the most important parameter regions of the two, which is the most direct way to avoid conflicts (visualization in the Appendix F.5). To verify this, we integrate the intersection set of the most important parameter regions of the safety and over-safety patches into the backbone model (denoted as “w/ Intersect. Patch” in Table 2). The slight changes of the resulting model in both safety enhancement and over-safety mitigation highlights the importance of addressing the conflicts in safety patching.
Table 3: Results on the safety, over-safety and utility benchmarks for continual PSA with LLaMA-2-Chat-7B backbone. All results are the average over the three time steps.
[…]
5.3 SafePatching for Continual Post Safety Alignment
We evaluate the effectiveness of SafePatching in a more challenging and practical continual PSA setting. Specifically, we select three categories of harmful data from Beavertails benchmark, drug abuse, animal abuse, and misinformation and law, to simulate users’ PSA needs at three different time steps. Continual PSA requires the model to use only the current harmful data for training at each time step, achieving the current safety improvement goals without compromising the previous ones, and ensuring that over-safety is mitigated while maintaining utility throughout the process. We present the average performance of different PSA methods at the three time steps in Table 3 and depict the fine-grained changes in performance over time in Figure 5. Overall, SafePatching effectively achieves safety improvement, over-safety mitigation, and utility maintenance compared to the baselines, both during and after the continual PSA process.
@@ 6 Conclusion
In this paper, we introduce SafePatching, a novel framework designed for the comprehensive and efficient post-safety alignment (PSA) of aligned large language models (LLMs). SafePatching develops two distinct types of safety patches for safety enhancement and over-safety mitigation, utilizing the same set of harmful safety-alignment data with different gradient optimization techniques. Then through controllable patching, they are seamlessly integrated into the target LLM backbone without compromising, and potentially even increasing, its utility. Experimental results demonstrate that SafePatching outperforms baseline methods in comprehensive and effective PSA, with lower training and inference costs, underscoring its efficiency. Furthermore, SafePatching remains effective in practical continual PSA settings.
Appendix A Limitations
Despite our SafePatching achieving comprehensive and efficient post-safety alignment (PSA), there are still limitations and future directions worth exploring:
  • Limited Scale of Backbones: Due to limited computational resources, our experiments are not conducted on larger-scale base models (larger than 13B). This remains a valuable direction to apply SafePatching on larger LLMs.

### v2

@@ 4 Experiments
4.1 Experimental Setup
Models
We select 4 representative aligned LLMs: LLaMA-2-7B-Chat (Touvron et al., 2023b), LLaMA-3-8B-Instruct (AI@Meta, 2024), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023) and Gemma-1.1-7B-it (Team et al., 2024), to fully validate the effectiveness and scalability of our SafePatching in achieving comprehensive PSA.
Safety Benchmark
We assess the safety under jailbreak attacks, using both Seen and Unseen harmful samples from AdvBench (Zou et al., 2023). In the Seen scenario, we directly simulate PSA as a “post hoc” defense against jailbreak attacks, training and testing PSA methods on the same jailbreak dataset split. No jailbreak templates are involved during the training process. Meanwhile, the scenario Unseen evaluates the generalizability of the PSA method on data that was not encountered during training. We consider four state-of-the-art jailbreak attacks that cover different categories, including ICA (Wei et al., 2023b), GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024c) and PAIR (Chao et al., 2024). Detailed description and setup can be found in Appendix C.1 and Appendix D.
We adopt Attack Success Rate (ASR) as the evaluation metric and it is calculated by a Longformer-based classifier (Beltagy et al., 2020) provided by Wang et al. (2023). It shows a performance identical to that of GPT-4 and human annotators to judge whether a response is harmful or not. Please refer to Appendix E for more details on the classifier.
Over-Safety Benchmark
The evaluation for over-safety mitigation is performed on XSTest (Röttger et al., 2024) and OKTest (Shi et al., 2024). As suggested by Röttger et al. (2024), we hire a human annotator to manually evaluate refusal rate. Please refer to Appendix F for more details.
Utility Benchmark Following Touvron et al. (2023b), we conduct utility evaluations across four crucial dimensions: (1) World Knowledge: MMLU (5-shot) (Hendrycks et al., 2020); (2) Reasoning: 0-shot for HellaSwag (Zellers et al., 2019), ARC-challeng (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021) and BBH (3-shot) Suzgun et al. (2023); (3) MATH: GSM8K (8-shot) (Cobbe et al., 2021) and MATH (4-shot) (Hendrycks et al., 2021) and (4) Multi-Turn Instruction-Following: MT-bench (Zheng et al., 2024b).
@@ 4.2 Baselines and Comparison Methods
We evaluate SafePatching against the following state-of-the-art PSA baselines:
Safety Enhancement: (1) GA (Yao et al., 2023), (2) GA + Mismatch (Yao et al., 2023), (3) $\text{RESTA}_{d}$ (Bhardwaj et al., 2024) (4) NPO (Zhang et al., 2024a) (5) SafeDecoding (Xu et al., 2024) and (6) ROSE (Zhong et al., 2024).
Over-Safety Mitigation: (1) Self-CD (Shi et al., 2024), (2) SCAN (Cao et al., 2024) and (3) Surgery (Wang et al., 2024).
To verify the advantages of the controllable patching in conflict mitigation, we also compare SafePatching with state-of-the-art model merging methods: (1) Average Merging (Wortsman et al., 2022), (2) Task Arithmetic (Ilharco et al., 2022a), (3) Fisher Merging (Matena and Raffel, 2022) and (4) TIES-Merging (Yadav et al., 2023).
Please refer to Appendix C for the detailed description of the baseline methods.
@@ 4.3 Implementation Details
Our experiments are implemented with PyTorch (Paszke et al., 2019) on 4 NVIDIA Tesla A100 using DeepSpeed (Rasley et al., 2020). A comprehensive list of hyper-parameters is provided in Table 5 within Appendix D. For further implementation details, please see Appendix D.
@@
Seen AVG. Unseen AVG. XSTest OKTest AVG. MT-Bench
LLaMA-2-7B-Chat 24.00 21.00 8.00 4.67 40.35 6.01
Average Merging (Wortsman et al., 2022) 13.50 14.00 9.67 10.00 40.35 6.03
Task Arithmetic (Ilharco et al., 2022b) 7.50 7.50 7.33 8.40 40.34 6.08
Fisher Merging (Matena and Raffel, 2022) 72.50 76.00 0 0 40.20 6.15
TIES-Merging (Yadav et al., 2023) 12.00 11.00 7.67 6.00 40.43 6.09
w/ Intersect. Patch 10.75 10.25 6.67 4.80 39.79 6.03
w/o Rand. Retention 8.00 8.50 8.33 6.40 40.41 5.78
w/ Safety Patch 11.25 7.00 10.67 12.80 39.27 5.88
w/ Over-Safety Patch 81.00 86.00 0 0 39.28 5.76
SafePatching 7.25 7.50 3.33 2.33 40.42 6.14
Table 2: The overall ablation results to verify the effeficay of different components in SafePatching. We compare it with the SOTA model merging methods. LLaMA-2-7B-Chat is the backbone. Results highlighted in green indicate an improvement or performance comparable to the original backbone, while those highlighted in red signal a decline in performance relative to the original backbone.
@@ 4.4 Overall Results
Table 1 demonstrates the performance comparison of SafePatching and baselines based on LLaMA-2-7B-Chat. Please refer to Appendix G.1 for more results on LLaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.1 and Gemma-1.1-7B-it.
SafePatching strikes a good balance between safety enhancement, over-safety mitigation and utility preservation across different backbones. Based on the results from all backbones in Table 1, as well as Tables 8, 9 and 10 in the Appendix G.1, we derive two key insights: (1) Striking a balance between safety and over-safety is extremely difficult, and none of the baseline methods manage to achieve this balance. Optimizing one aspect inevitably causes a substantial decline in the other. In contrast, our SafePatching optimizes both simultaneously, delivering the best balance compared to the baselines. (2) Existing baseline methods fail to protect the backbone’s general utility from degradation, while our approach successfully preserves these capabilities, particularly for multi-turn instruction-following. In conclusion, our SafePatching succeed to address the above challenging trilemma, comprehensively achieving all three goals of PSA.
@@
SafePatching could even enhance utility. Surprisingly, SafePatching outperforms the original LLaMA-2-7B-Chat and Gemma-1.1-7B-it on the MT-bench. We explore the reasons behind this improvement by analyzing GPT-4o’s evaluation of the generated responses. We discover that responses from SafePatching more frequently address the user’s input directly, unlike the original model, which may criticize the phrasing of user queries (all input prompts in MT-bench are safe). Detailed analysis can be found in Appendix H.3.
Figure 3: Visualization of the difference set of the most important parameters from the safety patch (left) and over-safety patch. The backbone is LLaMA-2-7B-Chat.
5 Analysis
5.1 Ablation Study
We conduct ablation experiments to analyze the efficacy of each component of SafePatching.
Effect of Safety and Over-Safety Patches
Whether incorporating safety patch or over-safety patch alone (denoted as “w/ Safety Patch” and “w/ Over-Safety Patch” in Table 2), the resulting model consistently excels in meeting the respective PSA objectives, indicating they are highly effective in achieving their specific goals. This also highlights the logic and effectiveness of using harmful data for gradient optimization to enhance safety and reduce over-safety independently.
Effect of Random Parameter Retention
When we remove the operation of random parameter retention (denoted as “w/o Rand. Retention” along with “w/ Safety Patch” and “w/ Over-Safety Patch” in Table 2), the utility of the resulting model decreases, revealing the conflict between parameters of these two patches and the backbone model.
Effect of Patching within the Difference Set
In Table 2, we compare SafePatching with the SOTA model merging methods. Note that they are also integrated with the same random retention techniques (Yu et al., 2024). Although TIES-Merging is the most effective one for mitigating conflicts, it struggles to find a balance between safety and over-safety. This issue arises from its lack of controlled merging in areas specifically related to safety and over-safety. More importantly, while these model merging methods succeed to increase the backbones’ utility, this enhancement has significantly compromised their safety. This trade-off is clearly unacceptable for delivering safe and reliable LLM services. We hope this trade-off serves as a reminder that future work on model merging should prioritize both improving utility and maintaining safety.
To further verify the impact of difference set, we integrate the intersection set of the most important parameter regions of the safety and over-safety patches into the backbone model (denoted as “w/ Intersect. Patch” in Table 2). The slight changes of the resulting model in both safety enhancement and over-safety mitigation highlights the importance of addressing the conflicts in safety patching.
5.2 Distribution of Important Parameters
The distribution of the most important parameters for safety and over-safety exhibits different patterns.
We show the difference set between the most important parameter regions for safety patch and over-safety patch in Figure 3. Our analysis reveals two main findings. (1) The important parameters related to over-safety are mostly concentrated in the lower transformer layers of the backbone, whereas those related to safety are relatively more evenly distributed, primarily in the middle layers of the model. This directly demonstrates that our method can alleviate the conflict between safety and over-safety to patch them in the different regions. (2) The important parameters for safety are predominantly located in the Feed Forward (FFN) layers, while those for over-safety are more concentrated in the attention layers. This supports the findings of Shi et al. (2024), which suggest that the issue of over-safety in current aligned LLMs is due to the their tendency to overly focus on sensitive words via the attention operation in the input prompt rather than fully understanding the true intention behind the user’s prompt.
5.3 Hyper-Parameter Analysis
The hyper-parameters in SafePatching exhibit consistent robustness across various backbones and the process for adjustment is very quick and efficient.
We conduct an analysis of hyper-parameters, including overall retention rate $p$, top rate $a$ and $b$, and scale weight $\alpha$ and $\beta$, across all backbone models. We can draw two conclusions: (1) Random retention rate $p$ and the top rate $a$ and $b$ are robust across different backbones. And top rate also exhibits robustness within each single backbone. (2) The only model-specific hyper-parameter we need to adjust is the scale weight $\alpha$ and $\beta$. And larger $\alpha$ would enhance safety performance but increase over-safety. Conversely, a larger $\beta$ has the opposite effect. The detailed results and analysis can be found in Appendix G.3.
@@ 5.3 Hyper-Parameter Analysis
Table 3: Results on safety, over-safety and utility benchmarks for continual PSA with LLaMA-2-Chat-7B backbone. Results are averaged over the three time steps.
5.4 Continual Post Safety Alignment
We evaluate SafePatching under the more challenging and practical continual PSA setting. Specifically, we select three categories of harmful data from Beavertails (Ji et al., 2023) benchmark, drug abuse, animal abuse, and misinformation and law, to simulate users’ different PSA needs at three time steps. Continual PSA requires the model to use only the current harmful data for training at each time step, achieving the current safety improvement goals without compromising the previous ones, and ensuring that over-safety is mitigated while maintaining utility throughout the process. Detailed setting can be found in Appendix G.4. We present the average performance of different PSA methods at the three time steps in Table 3 and depict the fine-grained changes in performance over time in Figure 9 in Appendix G.4. Overall, SafePatching still effectively achieves three PSA goals.
@@ 6 Conclusion
In this paper, we introduce SafePatching, a novel framework designed for the comprehensive post-safety alignment (PSA) of aligned large language models (LLMs). SafePatching develops two distinct types of safety patches for safety enhancement and over-safety mitigation, utilizing the same set of harmful safety-alignment data with different gradient optimization techniques. Then through controllable patching, they are seamlessly integrated into the target LLM backbone without compromising, and potentially even increasing, its utility. Experimental results on four representative aligned LLMs demonstrate that SafePatching outperforms baseline methods in comprehensive and effective PSA, underscoring its scalability. Furthermore, SafePatching remains effective in practical continual PSA settings.
7 Limitations
Despite our SafePatching achieving comprehensive post-safety alignment (PSA), there are still limitations and future directions worth exploring:
  • Limited Scale of Backbones: Due to limited computational resources, our experiments are not conducted on larger-scale base models (larger than 8B). This remains a valuable direction to apply SafePatching on larger LLMs.
@@ 7 Limitations
8 Ethical Considerations
In our pursuit to achieve post safety alignment in current aligned large language models (LLMs), it is imperative to underscore that, SafePatching could make these models safer to use, thereby increasing public trust in AI. This would encourage wider adoption of LLMs in various applications, from customer service to educational tools, knowing that the risk of harmful or inappropriate responses is minimized. And SafePatching’s efficiency in post safety alignment makes it easier to implement safety measures across different models and scenarios, potentially leading to a standardized approach to AI safety. This can streamline regulatory compliance and foster innovation while ensuring ethical AI practices.
While our SafePatching is designed to balance safety and usability by allowing benign responses to sensitive keywords, we recognize the potential risk of misuse by bad actors. As with any powerful technology, there is a risk that enhanced LLMs could also be misused. Even with improved safety mechanisms, malicious actors might find ways to exploit these models for disinformation, fraud, or other harmful purposes. Ensuring robust safeguards and monitoring systems will be crucial. To this end, we emphasize research-only usage, ensuring SafePatching is used solely for advancing safety research.
Appendix A Difference Between Over-Safety and Utility
Although inputs in both over-safety and utility evaluations are safe, they are conceptually different.
Over-safety pertains to adversarial testing with inputs that contain sensitive words but are otherwise harmless, measuring whether the model becomes overly sensitive and rejects benign inputs. Here are two examples from the XSTest (Röttger et al., 2024) dataset.
  • How do I make my wife explode with laughter?
  • How can I build a weapon in Minecraft?
The first example employs metaphorical language to construct a sensitive expression, while the second involves risky behavior in a controlled environment. However, both are essentially safe prompts, crafted with adversarial intent to challenge whether the model will refuse to respond.
Utility refers to the model’s helpfulness in practical applications, such as world knoledge, reading comprehension, reasoning and instruction-following capabilities. It primarily measures the model’s ability to solve specific tasks. The two, combined with the safety evaluation, jointly provide a comprehensive assessment of the current performance of the PSA methods.